# Casimir-Lifshitz interaction between bodies integrated in a micro/nanoelectromechanical quantum damped oscillator

### v1

@@ Abstract
A theory is proposed for the component of the Casimir-like force that arises between bodies embedded in a macroscopic quantum damped oscillator. When the oscillator’s parameters depend on the distance between the bodies, the oscillator-induced Casimir-like force is generally determined by a broad spectral range extending to high frequencies, limited by the frequency dispersion of the damping function. Here it is shown that there is a large class of systems in which the low-frequency range dominates the forces. This allows for the use of the Ohmic approximation, which is crucial for extending the theory to the lumped element description of fluctuation-induced forces in electrical circuits. Estimates of the circuit-induced Casimir-Lifshitz force suggest that under certain conditions it can be identified experimentally due to its dependence on various circuit elements.
@@
Introduction.Casimir forces originate from quantum fluctuations of the electromagnetic vacuum in the presence of boundaries or other inhomogeneities [1, 2]. Extending the theory to both quantum and thermal electromagnetic fluctuations in inhomogeneous equilibrium condensed matter systems, led to a general theory of Casimir and van der Waals (or Casimir-Lifshitz) forces between atoms, molecules, and macroscopic bodies [3, 4, 5] (see also reviews [6, 7, 8, 9, 10, 11]).
@@
Advances in micro- and nanophysics technologies, particularly in the development of mechanical transducers, atomic force microscopes, torsion pendulums, and oscillators, enabled detailed quantitative studies of Casimir-Lifshitz forces (see, for example, [12, 13, 14, 15, 10, 16, 17, 18, 19, 11, 20, 21, 22, 23, 24, 25, 26, 27] and references therein). These advances also drew significant attention to the influence of Casimir-Lifshitz forces on the performance of micro- and nanoelectromechanical devices, whose components often operate in close proximity [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49].
@@
This paper presents a theory of fluctuation-induced interaction that arises between bodies embedded in quantum damped oscillators as encountered in micro- or nanoelectromechanical devices. The Casimir-like force considered here is driven by the oscillator through a parametric dependence of its frequency $\Omega$ and/or its damping function $\gamma(\omega)$ on the distance between the bodies. In electrical circuits, a similar Casimir-Lifshitz force component emerges from electromagnetic eigenmodes shaped by distance-dependent lumped elements. This component can, under certain conditions, be detected experimentally, typically as a correction to the main Casimir-Lifshitz force. It can be distinguished by its dependence on various circuit elements.
@@
The position and momentum correlation functions, the partition function, and some related thermodynamic quantities of the quantum damped oscillator were studied in detail, mostly using the Zwanzig-Caldeira-Leggett model and assuming a bilinear coupling with the environment [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. These results can, in particular, be represented as sums or products over the Matsubara frequencies $\omega_{n}=2\pi nT/\hbar$.
@@
The quantities in question can be divided into two groups, depending on the spectral range of fluctuations that dominate them. The first group, for example, includes the mean-square momentum fluctuation and the ground state energy. The corresponding sums diverge at high frequencies in the Ohmic approximation, because their convergence is ensured only by the frequency dispersion of the damping function $\gamma(i\omega_{n})$. This dispersion is determined by the spectral density of the environment [51, 53, 60], which decays rapidly with $\omega_{n}$ at high frequencies, thereby defining a characteristic frequency range $\omega_{n}\lesssim\omega_{c}$ that dominates the quantities in the first group.
@@
By contrast, the quantities of the second group are primarily governed by fluctuations in the spectral region $\omega\lesssim\tilde{\Omega}$, where $\tilde{\Omega}\sim\Omega$ for small or moderate $\gamma$. When $\omega_{c}\gg\tilde{\Omega}$, that is applicable to many problems and is assumed below, the condition $\omega\lesssim\tilde{\Omega}$ is favorable for applying the Ohmic approximation or the lumped element approach. The mean squared position fluctuations of the damped oscillator and its specific heat are representative of this second group. Unlike the ground state energy, the specific heat of the quantum damped oscillator belongs to the second group since its behavior is determined by the temperature derivatives of the thermal part of the energy [58, 59, 61].
@@
The oscillator-induced Casimir-like force provides an instructive example in this regard. Representing the distance derivative of the full free energy of the damped oscillator, the force belongs to the first group if the damping function depends on distance. In this case, the Casimir-like force includes a high-frequency contribution limited only by the damping function, whose frequency dispersion must be specified. The corresponding result is obtained below using the Drude model.
The role of frequency dependent parameters is crucial for extending the theory of oscillator-induced Casimir-like forces to electrical circuits with a lumped element description, when the eigenmodes have frequency-independent damping functions. The divergences that arise in this case present a major challenge to achieving this goal. This paper identifies possible solutions to the problem that occur under certain conditions.
@@
Here we show that the low-frequency range dominates the forces only when the resonant frequency $\Omega$ depends on distance, while the damping function $\gamma$ does not. This makes a lumped element description of fluctuation-induced forces in electrical circuits possible. A similar description also applies to another situation, where either $\gamma$, or both $\Omega$ and $\gamma$, depend on distance, but only $\Omega$ depends on an additional parameter $\varkappa$. In this case, the force difference $f_{\Omega_{1}}-f_{\Omega_{2}}$, with $\Omega_{1,2}=\Omega(\varkappa_{1,2})$, is governed by the low-frequency range.
@@
Oscillator-induced Casimir-like forces.Consider a quantum macroscopic damped oscillator, assuming that its resonant frequency $\Omega$ and the damping function $\gamma$ depend on the distance $d$ between the constituent interacting bodies. These quantities enter the quantum Langevin equation for the oscillator’s coordinate $M\ddot{\hat{Q}}+M\int_{0}^{t}d\tau\gamma(t-\tau)\dot{\hat{Q}}(\tau)+M\Omega^{2}\hat{Q}=\hat{\xi}(t)$.
@@
It follows from the expression for the damped oscillator’s free energy $F(d)=-T\log Z(d)$, known in the Zwanzig-Caldeira-Leggett model together with the partition function $Z(d)$ [53, 55, 60, 63], that the interaction force $f=-\frac{\partial F}{\partial d}$ takes the form  [^1] See Supplemental Material for details of the calculations of the oscillator-induced Casimir-like force under various conditions, as well as of the Casimir-Lifshitz force induced by the RLC circuits in series and in parallel.
@@
Here $\gamma(\omega)=\int_{0}^{\infty}\gamma(t)e^{i\omega t}dt$, the summation is taken over the Matsubara frequencies, and the prime at the summation sign indicates that the term with $n=0$ is taken with half weight.
In (1), each of the contributions to the interaction force, $f_{\Omega}$ or $f_{\gamma}$, describes repulsion when $\Omega$ or $\gamma(i\omega_{n})$, respectively, decreases with distance, and attraction when it increases.
The term $f_{\gamma}\propto\frac{\partial\gamma(i\omega_{n},d)}{\partial d}$ in (1) belongs to the first group because it is sensitive to and limited by the frequency dependence of $\gamma(i\omega_{n})$, and logarithmically diverges in the Ohmic approximation. By contrast, the term $f_{\Omega}\propto\frac{\partial\Omega(d)}{\partial d}$ in (1) is dominated by a comparatively low frequency range, which is $\omega_{n}\lesssim\Omega$ for small and moderate $\gamma$. Considering that $\omega_{c}\gg\Omega$, the force component $f_{\Omega}$ can be described in the Ohmic approximation and thus belongs to the second group.
Focusing on the oscillator- or circuit-induced forces shaped mostly by low-frequency fluctuations, we assume first that only $\Omega$ depends on $d$, while $\gamma(\omega)$ remains fixed. (Examples will be given below.) Then one obtains from (1) the following expression for the total fluctuation force $f(d)=f_{\Omega}(d)$ in the Ohmic approximation:
@@
where the quantities $i\omega_{1,2}=\frac{\gamma}{2}\pm i\sqrt{\Omega^{2}-\frac{\gamma^{2}}{4}}$ enter the arguments of the digamma functions $\psi(x)$.
For weak dissipation $\gamma^{2}\ll 4\Omega^{2}$, the force (2) is given by the derivative of the free energy of the quantum harmonic oscillator $f=-\frac{1}{2}\coth\frac{\hbar\Omega}{2T}\frac{\partial\Omega}{\partial d}$. At low temperatures $T\ll\hbar\Omega$, this reduces to the contribution from the zero-point energy $f_{\Omega}=-\frac{\hbar}{2}\frac{\partial\Omega}{\partial d}$. At high temperatures $\hbar\Omega\ll T$, the first term in (2) dominates the force.
If $\gamma$ is not negligibly small, the dominating term that follows from (2) in the low temperature limit is
@@
In the limit of a strong dissipation, $4\Omega^{2}\ll\gamma^{2}$, the frequency $i\omega_{2}\approx\frac{\Omega^{2}}{\gamma}$ becomes anomalously small with increasing $\gamma$, while another one linearly increases $i\omega_{1}\approx\gamma$. The low temperature range is $T\ll\frac{\Omega^{2}}{\gamma}$ in this case, while the value of $\gamma$ is confined by the condition $\gamma\ll\omega_{c}$.
@@
If $\gamma(\omega,d)$ depends on distance $d$, possibly together with $\Omega$, then an additional term $f_{\gamma}\propto\frac{\partial\gamma}{\partial d}$ appears in the expression of the total fluctuation force $f=f_{\Omega}+f_{\gamma}$, and a wide spectral range that includes the high frequences $\omega\lesssim\omega_{c}$ forms $f_{\gamma}$. Let $\Omega$, however, be dependent on an additional parameter $\varkappa$, while $\gamma(i\omega_{n},d)$ is not. One can see from (1) that the difference $f_{\gamma}(d,\varkappa_{1})-f_{\gamma}(d,\varkappa_{2})$ belongs to the second group, whereas the individual quantities $f_{\gamma}(d,\varkappa_{1,2})$ enter the first group.
Introducing the quantity $\tilde{f}(d,\varkappa)$, that satisfies the relation $\tilde{f}(d,\varkappa_{1})-\tilde{f}(d,\varkappa_{2})=f(d,\varkappa_{1})-f(d,\varkappa_{2})$ and retains only those terms from $f(d,\varkappa)$, which generally do not cancel out in this difference, we obtain within the Ohmic approximation
@@
For describing the total oscillator-induced Casimir-like force in the presence of distance-dependent $\gamma$, the frequency dependence of $\gamma$ must be specified in (1). Consider here the Drude-like model $\gamma(\omega)=\frac{\gamma_{0}\omega_{D}}{\omega_{D}-i\omega}$, which is commonly used in studying the damped oscillator properties [52, 53, 55], and assume that generally all three model parameters, $\Omega$, $\gamma_{0}$ and $\omega_{D}$, depend on $d$. A pronounced frequency dependence of $\gamma(i\omega,d)$ shows up in the model in the range $\omega\gtrsim\omega_{D}$, so that $\omega_{D}$ plays in (1) the role of an effective cutoff frequency $\omega_{c}$.
In the Drude model each term involving $\frac{\partial\gamma_{0}}{\partial d}$ or $\frac{\partial\omega_{D}}{\partial d}$ describes repulsion when the parameter $\gamma_{0}$ or $\omega_{D}$ decreases with $d$, and attraction when it increases.
When the frequency dispersion of $\gamma$ is disregarded, the damped oscillator has two complex eigenfrequencies $\omega_{1,2}$, as used in (2) and (4), whereas in the Drude model it has three, $\omega_{1,2,3}$. If $\omega_{D}\gg\Omega,\gamma_{0}$, as is assumed here, these frequencies can be readily obtained using an expansion in powers of the small parameters $\frac{\Omega}{\omega_{D}}$ and $\frac{\gamma_{0}}{\omega_{D}}$ [52]. To first order, $i\omega_{1,2}=\frac{\gamma_{0}}{2}\pm i\sqrt{\Omega^{2}-\frac{\gamma_{0}^{2}}{4}}$ and $i\omega_{3}=\omega_{D}-\gamma_{0}\gg|\omega_{1,2}|$. In this case, the total fluctuation force is $f(d)=\tilde{f}(d)+\Delta f_{\gamma}(d)$ [64]. The first term is defined in (4), where $\gamma_{0}$ must be substituted for $\gamma$, and the second term contains the third eigenfrequency and the parameter $\omega_{D}$:
@@
Three contributions to the interaction force in (6) arise from the distance dependence of $\Omega$, $\gamma_{0}$ and $\omega_{D}$ and can generally have the opposite signs. The large argument under the logarithm sign in (6) reflects the logarithmic divergence of the sum in (1), which appears when the frequency dispersion of $\gamma$ is neglected.

### v2

# Casimir-Lifshitz interaction between bodies integrated in a microelectromechanical/nanoelectromechanical quantum damped oscillator

@@ Abstract
A theory is proposed for the component of the Casimir-like force that arises between bodies embedded in a macroscopic quantum damped oscillator. When the oscillator’s parameters depend on the distance between the bodies, the oscillator-induced Casimir-like force is generally determined by a broad spectral range extending to high frequencies, limited by the frequency dispersion of the damping function. Here it is shown that there exists a large class of systems in which the low-frequency range dominates the forces. This allows one to use the Ohmic approximation, which is crucial for extending the theory to a lumped element description of fluctuation-induced forces in electrical circuits. Estimates of the circuit-induced Casimir-Lifshitz force suggest that under certain conditions it can be identified experimentally due to its dependence on various circuit elements.
I Introduction
@@
Casimir forces originate from quantum fluctuations of the electromagnetic vacuum in the presence of boundaries or other inhomogeneities [1, 2]. Extending the theory to both quantum and thermal electromagnetic fluctuations in inhomogeneous equilibrium condensed matter systems, resulted in a general theory of Casimir and van der Waals (or Casimir-Lifshitz) forces between atoms, molecules, and macroscopic bodies [3, 4, 5] (see also reviews [6, 7, 8, 9, 10, 11]).
@@
Advances in micro- and nanophysics technologies, particularly in mechanical transducers, atomic force microscopes, torsion pendulums, and oscillators, enabled detailed quantitative experimental studies of Casimir-Lifshitz forces (see, for example, [12, 13, 14, 15, 10, 16, 17, 18, 19, 11, 20, 21, 22, 23, 24, 25, 26, 27] and references therein). These advances also drew significant attention to the influence of Casimir-Lifshitz forces on the performance of micro- and nanoelectromechanical devices, whose components often operate in close proximity [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49].
@@
This paper proposes a theory of fluctuation-induced interactions between bodies embedded in quantum damped oscillators as encountered in micro- or nanoelectromechanical devices. The Casimir-like force considered here is driven by the oscillator through a parametric dependence of its frequency $\Omega$ and/or its damping function $\gamma(\omega)$ on the distance between the bodies. In electrical circuits, a similar Casimir-Lifshitz force component emerges from electromagnetic eigenmodes shaped by distance-dependent lumped elements. This component can, under certain conditions, be detected experimentally, typically as a correction to the main Casimir-Lifshitz force. It can be distinguished by its dependence on various circuit elements.
@@
The conventional microscopic approach to quantum effects in dissipative systems is based on the Hamiltonian of the entire conservative system, which includes the original system, its environment (thermal bath), and their interaction terms. The position and momentum correlation functions, the partition function, and some related thermodynamic quantities of the quantum damped oscillator have been studied in detail, mostly within the Zwanzig-Caldeira-Leggett model with bilinear coupling to the environment [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. These quantities can, in particular, be represented as sums or products over the Matsubara frequencies $\omega_{n}=2\pi nT/\hbar$.
@@
The quantities in question can be divided into two groups, depending on the spectral range of fluctuations that dominate them. The first group, for example, includes the mean-square momentum fluctuation and the ground state energy. The corresponding sums diverge at high frequencies in the Ohmic approximation. A constant damping function over the entire spectral range implies the presence of a thermal bath with an infinite number of degrees of freedom, which links these divergences to those encountered in the quantum-field Casimir effect. Convergence then is ensured only by the frequency dispersion of the damping function $\gamma(i\omega_{n})$, which is determined by the spectral density of the environment [51, 53, 60] decaying rapidly with $\omega_{n}$ at high frequencies and defining a characteristic cutoff frequency range $\omega_{n}\lesssim\omega_{c}$ that governs the quantities in the first group.
@@
By contrast, the quantities of the second group are primarily determined by fluctuations in the spectral region $\omega\lesssim\tilde{\Omega}$, where $\tilde{\Omega}\sim\Omega$ for small or moderate $\gamma$. When $\omega_{c}\gg\tilde{\Omega}$, which is applicable to many problems and is assumed below, the condition $\omega\lesssim\tilde{\Omega}$ is favorable for applying the Ohmic approximation or a lumped element description. The mean squared position fluctuations of the damped oscillator and its specific heat are representative of the second group. Unlike the ground state energy, the specific heat belongs to the second group because it is determined by temperature derivatives of the thermal part of the energy [58, 59, 61].
@@
The oscillator-induced Casimir-like force provides an instructive example in this regard. Representing the distance derivative of the full free energy of the damped oscillator, the force, as shown below, belongs to the first group if the damping function depends on distance. In that case, the Casimir-like force includes a high-frequency contribution limited only by the frequency dependence of $\gamma(i\omega_{n})$, which must then be specified explicitly. The corresponding results will be obtained using the Drude model.
The role of frequency dependent parameters is crucial for extending the theory of oscillator-induced Casimir-like forces to electrical circuits with a lumped element description. Divergences that generally arise when describing forces in lumped electrical circuits present a major theoretical challenge. Treating a quantum damped oscillator as a prototype for the electric-circuit eigenmodes, this paper identifies simple and general conditions under which finite (free of divergences) results for the corresponding contribution to the force can be obtained within the lumped-circuit description  [^1] This paper considers only lumped circuits composed of lumped elements. In transmission-line circuits, effects of propagating waves along the circuit wires can substantially affect the results [90, 89].. The results are derived from the solution for a damped harmonic oscillator within the Zwanzig-Caldeira-Leggett model.
@@
We show that the low-frequency range dominates the force only when the resonant frequency $\Omega$ depends on distance, whereas the damping function $\gamma$ does not. This enables the use of a lumped-element description of fluctuation-induced forces in electrical circuits. A similar description applies to cases where either $\gamma$, or both $\Omega$ and $\gamma$, depend on distance, but only $\Omega$ depends on an additional parameter $\varkappa$. In that case, the force difference $f_{\Omega_{1}}-f_{\Omega_{2}}$, with $\Omega_{1,2}=\Omega(\varkappa_{1,2})$, is governed by the low-frequency range. Explicit expressions for these forces are given below for two particular cases.
II Oscillator-induced Casimir-like forces
@@
Consider a quantum macroscopic damped oscillator, whose resonant frequency $\Omega$ and damping function $\gamma$ depend on the distance $d$ between the constituent interacting bodies. These quantities enter the quantum Langevin equation for the oscillator’s coordinate $M\ddot{\hat{Q}}+M\int_{0}^{t}d\tau\gamma(t-\tau)\dot{\hat{Q}}(\tau)+M\Omega^{2}\hat{Q}=\hat{\xi}(t)$.
@@
From the free energy $F(d)=-T\log Z(d)$ of the damped oscillator in the Zwanzig-Caldeira-Leggett model, which comes together with the partition function $Z(d)$ [53, 55, 60, 63], it follows that the interaction force $f=-\tfrac{\partial F}{\partial d}$ can be written as  [^2] See Supplemental Material at http://link.aps.org/ supplemental/10.1103/b6kg-nykc for details of the calculations of the oscillator-induced Casimir-like force under various conditions, as well as of the Casimir-Lifshitz force induced by the series and parallel RLC circuits. The Supplemental Material also contains Refs. [91, 92, 93].
@@
Here $\gamma(\omega)=\int_{0}^{\infty}\gamma(t)e^{i\omega t}dt$, the summation is over Matsubara frequencies, and the prime at the summation sign indicates that the term with $n=0$ is taken with half weight.
In (1), each of the contributions, $f_{\Omega}$ or $f_{\gamma}$, describes repulsion, when $\Omega$ or $\gamma(i\omega_{n})$ respectively decreases with distance, and describes attraction when it increases.
The term $f_{\gamma}\propto\tfrac{\partial\gamma(i\omega_{n},d)}{\partial d}$ belongs to the first group because it is sensitive to and limited by the frequency dependence of $\gamma(i\omega_{n})$, and logarithmically diverges in the Ohmic approximation. By contrast, the term $f_{\Omega}\propto\tfrac{\partial\Omega(d)}{\partial d}$ is dominated by a comparatively low frequency range, $\omega_{n}\lesssim\Omega$ for small and moderate $\gamma$. For $\omega_{c}\gg\Omega$, the force component $f_{\Omega}$ can be treated in the Ohmic approximation and thus belongs to the second group.
Focusing on oscillator- or circuit-induced forces shaped mostly by low-frequency fluctuations, we first assume that only $\Omega$ depends on $d$, while $\gamma(\omega)$ remains fixed (examples to follow). Then one obtains from (1) the following expression for the total fluctuation force $f(d)=f_{\Omega}(d)$ in the Ohmic approximation:
@@
where the quantities $i\omega_{1,2}=\tfrac{\gamma}{2}\pm i\sqrt{\Omega^{2}-\tfrac{\gamma^{2}}{4}}$ enter the arguments of the digamma functions $\psi(x)$.
For weak dissipation $\gamma^{2}\ll 4\Omega^{2}$, the force (2) is given by the derivative of the free energy of the quantum harmonic oscillator $f=-\tfrac{1}{2}\coth\tfrac{\hbar\Omega}{2T}\tfrac{\partial\Omega}{\partial d}$. At low temperatures $T\ll\hbar\Omega$, this reduces to the zero-point energy contribution $f_{\Omega}=-\tfrac{\hbar}{2}\tfrac{\partial\Omega}{\partial d}$. At high temperatures $\hbar\Omega\ll T$, the first term in (2) dominates the force.
If $\gamma$ is not negligibly small, the dominant term that follows from (2) in the low temperature limit is
@@
In the limit of a strong dissipation, $4\Omega^{2}\ll\gamma^{2}$, the frequency $i\omega_{2}\approx\tfrac{\Omega^{2}}{\gamma}$ becomes anomalously small with increasing $\gamma$, while another one linearly increases $i\omega_{1}\approx\gamma$. The low temperature range is $T\ll\tfrac{\Omega^{2}}{\gamma}$ in this case, while the value of $\gamma$ is confined by the condition $\gamma\ll\omega_{c}$.
@@
When $\gamma(\omega,d)$ depends on distance $d$, possibly along with $\Omega$, then an additional term $f_{\gamma}\propto\tfrac{\partial\gamma}{\partial d}$ appears in the expression for the total fluctuation force $f=f_{\Omega}+f_{\gamma}$, and a wide spectral range that includes the high frequences $\omega\lesssim\omega_{c}$ forms $f_{\gamma}$. However, if $\Omega$ depends on an additional parameter $\varkappa$, while $\gamma(i\omega_{n},d)$ does not, then, as seen from from (1), the difference $f_{\gamma}(d,\varkappa_{1})-f_{\gamma}(d,\varkappa_{2})$ belongs to the second group, whereas the individual quantities $f_{\gamma}(d,\varkappa_{1,2})$ get included into the first group.
Introducing the quantity $\tilde{f}(d,\varkappa)$, that satisfies the relation $\tilde{f}(d,\varkappa_{1})-\tilde{f}(d,\varkappa_{2})=f(d,\varkappa_{1})-f(d,\varkappa_{2})$ and retains only those terms from $f(d,\varkappa)$, which generally are not canceled out in this difference, we obtain within the Ohmic approximation
@@
For describing the total oscillator-induced Casimir-like force in the presence of distance-dependent $\gamma$, the frequency dependence of $\gamma$ must be specified in (1). Consider here the Drude-like model $\gamma(\omega)=\tfrac{\gamma_{0}\omega_{D}}{\omega_{D}-i\omega}$, which is commonly used in studying the damped oscillator properties [52, 53, 55], and assume that generally all three model parameters, $\Omega$, $\gamma_{0}$ and $\omega_{D}$, depend on $d$. A pronounced frequency dependence of $\gamma(i\omega,d)$ shows up in the model in the range $\omega\gtrsim\omega_{D}$, so that in (1) $\omega_{D}$ plays the role of an effective cutoff frequency $\omega_{c}$.
In the Drude model each term involving $\tfrac{\partial\gamma_{0}}{\partial d}$ or $\tfrac{\partial\omega_{D}}{\partial d}$ describes repulsion, when the parameter $\gamma_{0}$ or $\omega_{D}$ respectively decreases with $d$, and attraction when it increases.
When the frequency dispersion of $\gamma$ is disregarded, the damped oscillator has two complex eigenfrequencies $\omega_{1,2}$, as used in (2) and (4), whereas in the Drude model it has three, $\omega_{1,2,3}$. If $\omega_{D}\gg\Omega,\gamma_{0}$, as is assumed here, these frequencies can be readily obtained using an expansion in powers of the small parameters $\tfrac{\Omega}{\omega_{D}}$ and $\tfrac{\gamma_{0}}{\omega_{D}}$ [52]. To the first order, $i\omega_{1,2}=\tfrac{\gamma_{0}}{2}\pm i\sqrt{\Omega^{2}-\tfrac{\gamma_{0}^{2}}{4}}$ and $i\omega_{3}=\omega_{D}-\gamma_{0}\gg|\omega_{1,2}|$. In this case, the total fluctuation force is $f(d)=\tilde{f}(d)+\Delta f_{\gamma}(d)$ [65]. The first term is defined in (4), where $\gamma_{0}$ must be substituted for $\gamma$, and the second term contains the third eigenfrequency as well as the parameter $\omega_{D}$:
@@
Three contributions to the interaction force in (6) arise from the distance dependence of $\Omega$, $\gamma_{0}$ and $\omega_{D}$ and, in general, can have the opposite signs. The large argument under the logarithm sign in (6) reflects the logarithmic divergence of the sum in (1), which appears when the frequency dispersion of $\gamma$ is neglected.
III Casimir-Lifshitz forces in circuits